The perimeter of a square is 32 feet. What is the area of the square?
Found 3 solutions by rfer, edjones, mananth:
Answer by rfer(16322)

32/4=8
8^2=64 sq ft

Answer by edjones(8007)

32/4=8 length of one side.
8^2=64 ft^2 Area.
.
Ed

Answer by mananth(16946)

perimeter = 32
let l be the length of the side of square
perimeter of square = 4l
4l=32
/4
4l/4 = 32/4
l =8 cm
..
Area = l^2
Area = 8*8
Area = 64 ft^2
..
m.ananth@hotmail.ca